THE INVERSION MOTION IN THE Ne - NH_{3} VAN DER WAALS DIMER STUDIED VIA MICROWAVE SPECTROSCOPY

LAURA E. DOWNIE^{a}, \underline{JULIE M. MICHAUD} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, CANADA.

The umbrella inversion motion of the ammonia molecule leads to a tunneling splitting of the ammonia energy levels. Transitions between rotational energy levels in different tunneling states lie in the microwave region and can be studied using a Fourier transform microwave spectrometer. Our study of Ne - NH_{3} investigates the effect of the neon atom on the NH_{3} inversion motion. Twelve transitions in total have been measured for the inversion motion in ^{20}Ne - ^{14}NH_{3} (and several minor isotopomers: ^{22}Ne - ^{14}NH_{3}, ^{20}Ne - ^{15}NH_{3}, ^{22}Ne - ^{15}NH_{3}). The gas sample contained 0.3% NH_{3} in 4-5 atm of Ne. Due to isotope enrichment of ^{22}Ne (9.25% naturally abundant) in the molecular expansion, the ^{22}Ne - NH_{3} signals had similar intensities to the ^{20}Ne - NH_{3} signals. To distinguish the isotopomers, an enriched sample of 3% ^{20}Ne (99.95%, Cambridge Isotopes) with 0.3% NH_{3} in a helium backing gas was used. The results from this enriched sample study show an anomalous isotope effect where the inversion transitions of the heavier ^{22}Ne isotopomers are observed at higher frequencies. This isotope effect is not observed with the Ne - ^{15}NH_{3} isotopomers, whose inversion transitions were measured at lower frequencies than the analogous Ne - ^{14}NH_{3} dimers.

aCurrent address: Department of Physics, University of New Brunswick, Fredericton, NB, CANADA